OpenStudy (anonymous):

1. Which of the following is the sum of the first 15 terms of the series 2, -12, 72, ...? 2. What is sum of the first 20 terms of the series -3, 0, 3, ...? 3. What is the sum of the first 30 terms of the series 9, 3, 1, ...?

OpenStudy (campbell_st):

well the series seems to be geometric... can you identify the common ratio...?

OpenStudy (campbell_st):

question 2. seems to be arithmetic.... can you identify the common difference Question 3. seems to be geometric... so you need a common ratio.

OpenStudy (anonymous):

Im just extremely confused... can you explain this to me at all?

OpenStudy (campbell_st):

on.... question 1 is a geometric series... you are multiplying the 1st term by a value to get the 2nd term... and then the 2nd term is multiplied by the same vale to get the 3rd term... etc... so the multiplying value is called the common ratio so 2 x ? = -12 -12 x ? = 72 any ideas of ?

OpenStudy (anonymous):

Yes, you have to multiply each term by -6

OpenStudy (campbell_st):

great you have the common ratio... so you need the formula for the sum of a geometric series \[S_{n} = \frac{a_{1}\times ( 1 - r^n)}{1 - r}\] in question 1. n = 15, a1 is the 1st term = 2 and r = -6 so substituting \[S_{15} = \frac{2 \times(1 - (-6)^{15})}{1 - (-6)}\] so now you can calculate the sum

OpenStudy (anonymous):

Would it be 134338567022?

OpenStudy (anonymous):

-2.5 x 1010 2.5 x 1010 -1.3 x 1011 1.3 x 1011 these are my options so Im confused

OpenStudy (campbell_st):

yep... that looks correct... now question 2... its an arithmetic sequence and the sum is \[S_{n} = \frac{n}{2} [ 2a_{1} + (n - 1) \times d]\] you know n = 20, a1 = -3..... you need to find d, which is what is added to each term.. substitute into the formula as calculate

OpenStudy (anonymous):

Wait can you look at my previous response to help me determine which is the answer for number 1? I got 510 for number 2

OpenStudy (campbell_st):

its the last one

OpenStudy (campbell_st):

thats correct... for 2 now look at 3... its geometric... so what do you multiply by...?

OpenStudy (anonymous):

3

OpenStudy (campbell_st):

no, you're dividing by 3... so how do you write that as a multiplication...?

OpenStudy (anonymous):

14.14?

OpenStudy (anonymous):

Now how would i solve this? Find the 40th term of the sequence 2, -12, 72

OpenStudy (campbell_st):

well this is the 1st question again, with n = 40

OpenStudy (anonymous):

But its not the sum, just finding the 40th term

OpenStudy (campbell_st):

for the 3rd question, the common ratio r = 1/3 a1 = 9 and n = 30 go back and use the geometric sum formula that's posted for the 1st question

OpenStudy (anonymous):

Wouldn't it come out to 14.14?

OpenStudy (campbell_st):

well I got 13.5

OpenStudy (anonymous):

Oh I see my mistake now in one of my calculations, Thanks! Could you just help me with th last one finding the 40th term?

OpenStudy (campbell_st):

well the 40th term question uses the same sequence as the 1st sequence just looking at things... all that is changed is n = 40 so redo question 1 with n = 40

OpenStudy (campbell_st):

oops ok to find the 4th term you need \[a_{n} = a_{1} \times r^{n - 1}\] so you know n, r and a1 just substitute and calculate

OpenStudy (campbell_st):

ooops 40th

OpenStudy (anonymous):

2^40*3^39?

OpenStudy (campbell_st):

well a1 = 2 and r = -6 and n = 40

OpenStudy (anonymous):

Im so confused!

OpenStudy (anonymous):

How would I show it in simplified terms?

OpenStudy (campbell_st):

\[a_{40} = 2 \times (-6)^{40 - 1}\] just calculate it out

OpenStudy (anonymous):

I did, but the number is so big, how can I show it in simplified form like the multiple choice options from first question?

OpenStudy (campbell_st):

well you index solution is correct

OpenStudy (anonymous):

???

OpenStudy (campbell_st):

you wrote 2^40 * 3^39

OpenStudy (anonymous):

Oh awesome thanks!